cookiemonster:

Functions f(x) and g(x) are shown below: f(x) = x^2 g(x) = x^2 + 8x + 16 In which direction and by how many units should f(x) be shifted to obtain g(x)? Left by 4 units Right by 4 units Left by 8 units Right by 8 units

cookiemonster:

@rootbeer003 @Tr3 @UsArmy3947 @demolisher1224 @lolokrat @psirockin2

cookiemonster:

@Shadow @sillybilly123 @563blackghost

sillybilly123:

complete the square on g(x) = x^2 + 8x + 16

cookiemonster:

@Shadow @ThisGirlPretty @Tr3

sillybilly123:

Completing the Square looks like this: \(g(x) = x^2 + 8x + 16 = (x + 4)^2 \) So: \(f(x) = g(x + 4)\)

cookiemonster:

right, so then what would the final answer be?

cookiemonster:

@sillybilly123